### v1

@@ Abstract
Accurate completion and denoising of roof height maps are crucial to reconstructing high-quality 3D buildings. Repairing sparse points can enhance low-cost sensor use and reduce UAV flight overlap. RoofDiffusion is a new end-to-end self-supervised diffusion technique for robustly completing, in particular difficult, roof height maps. RoofDiffusion leverages widely-available curated footprints and can so handle up to 99% point sparsity and 80% roof area occlusion (regional incompleteness). A variant, No-FP RoofDiffusion, simultaneously predicts building footprints and heights. Both quantitatively outperform state-of-the-art unguided depth completion and representative inpainting methods for Digital Elevation Models (DEM), on both a roof-specific benchmark and the BuildingNet dataset. Qualitative assessments show the effectiveness of RoofDiffusion for datasets with real-world scans including AHN3, Dales3D, and USGS 3DEP LiDAR. Tested with the leading City3D algorithm, preprocessing height maps with RoofDiffusion noticeably improves 3D building reconstruction. RoofDiffusion is complemented by a new dataset of 13k complex roof geometries, focusing on long-tail issues in remote sensing; a novel simulation of tree occlusion; and a wide variety of large-area roof cut-outs for data augmentation and benchmarking. Code and dataset: github.com/kylelo/RoofDiffusion
@@
Corrupted Point Cloud
Corrupted Height Map
Denoise 50%
Denoise 90%
Denoise 96%
Restored Height Map
Restored Point Cloud
@@
Figure 2: Types of corrupted roof height maps with real-world scan.
Refer to caption (a) Sparsity Refer to caption (b) Incomplete Refer to caption (c) Tree Noise
@@ 1 Introduction
Thanks to the efforts of contributors worldwide, OpenStreetMap [56] now increasingly provides vectorized high-precision footprints [29, 24]. Consequently, recent building reconstruction algorithms [29, 24, 2] have incorporated footprints into the reconstruction process. However, these approaches still struggle with sparsity, incompleteness and noise. Footprint-guided RoofDiffusionF addresses these long-tail challenges (Fig. 3).
@@
  • We unveil the PoznanRD with 13k Level of Detail (LoD) 2.2 [4] noise-free roof meshes and height maps.
    When treated with our noise and incompleteness algorithm, our complex roofs dataset can effectively generate ample training data to address “long tail” challenges, see Fig. 3.
@@ 4.2 Height Completion based on a Diffusion Model
Given $\boldsymbol{x}$, our goal is to generate a predicted image, $\hat{\boldsymbol{x}}_{0}$, that closely approximates the normalized ground truth image, $\boldsymbol{x}_{\text{gt}}$. We conceptualize this height completion task as a diffusion process. Drawing on the work of [25], our approach employs both a forward and a reverse Markov chain. The forward chain perturbs the ground truth data through noise injection, while the reverse chain utilizes a learnable model, conditioned on the input, to restore the data, see Fig. 4.
@@
The Reverse Process sequentially removes noise from the data to yield a complete, noise-free version. Specifically, we can draw a noisy image $\boldsymbol{x}_{T}$ from $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ at time $t=T$. Then use a reverse transition kernel to recursively eliminate noise until $t=0$, resulting in the restored height image, $\hat{\boldsymbol{x}}_{0}$. This transition kernel is parameterized by a learnable model $\boldsymbol{\epsilon}_{\theta}$. While $\hat{\boldsymbol{x}}_{0}$ conforms to $q(\boldsymbol{x}_{\text{gt}})$, it may not always correspond to the normalized input $\boldsymbol{x}$. In our case, the objective is to ensure that the repaired roofs seamlessly integrate with pre-existing structures. To ensure a close association between $\boldsymbol{x}$ and $\hat{\boldsymbol{x}}_{0}$, we condition $\boldsymbol{\epsilon}_{\theta}$ on $\boldsymbol{x}_{t}$, $\boldsymbol{x}$, and $t$, in the spirit of [70]. Each step of the reverse process can be formulated as
@@
where $\alpha_{t}$ and $\bar{\alpha}_{t}$ are hyperparameters for variance scheduling, and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. Furthermore, $\boldsymbol{\epsilon}_{\theta}$ depends on $\boldsymbol{m}\odot\boldsymbol{x}_{t}-\boldsymbol{m}^{\prime}$, the corrupt normalized data, $\boldsymbol{x}$, and $\bar{\alpha}_{t}$.
@@ 5 Datasets & Benchmarks
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Sparsity of real-world USGS 3DEP LiDAR Dataset [79, 78].
@@ 5.1 Tree Points
In analyzing height maps within a footprint, tree crowns may intrude, causing LiDAR to capture the tree canopy instead of the roof, see Fig. 1(c). Thus, we simulate tree intrusion by planting virtual trees around the roof. We collected a real-world database with 1k tree height maps from [10]. Small trees are randomly placed outside the footprint, with random rotations and height adjustments. A “max” operation is used to combine the roof $\boldsymbol{x}$ and tree height map $\boldsymbol{x}_{\text{tree}}$. Algorithm 3 in supplements. has the details.
@@ 6.1 Quantitative Evaluation of Height Completion
    (No-FP) Both sparsity and incompleteness on PoznanRD in Tab. 2.
@@
(No-FP) Both sparsity and incompleteness. We compare to and the state-of-the-art unguided depth completion methods, pNCNN [15] and CU-Net [85], that we selected for their exclusive use of depth data and hence close relation to height completion tasks. These algorithms were trained on our PoznanRD, using identical data augmentation settings, but no footprint information was used. We adhered to the default hyperparameter configurations specified for each method. To ensure a fair comparison, we also trained No-FP RoofDiffusion, a variant of RoofDiffusion that does not require footprints. We tested reconstruction accuracy on PoznanRD (s95 i30, s95 i60, s99 i30, and s99 i60) utilizing MAE and RMSE. The noise injection settings mirror the experiment in Tab. 1. Since predicting both footprint and roof simultaneously is a more difficult problem the sparsity and incompleteness selection in Tab. 2 is lower than in Tab. 1.
@@
(No-FP) Either sparsity or incompleteness. When tested on the BuildingNet [73] dataset, No-FP RoofDiffusion outperforms linear interpolation, CU-Net [85], and pNCNN [15] in scenarios with dominant sparsity (s98 i0, s99 i0) or incompleteness (s90 i60, s90 i80). While CU-Net’s smoother predictions reduce large errors and improve RMSE, notably at s90 i80, s98 i0, and s99 i0, CU-Net does not predict sharp height maps well, as indicated by a higher MAE.
@@ 6.3 Enhancement of 3D Reconstruction
Figure 9 illustrates that using RoofDiffusion as a preprocessing step significantly enhances the reconstruction quality in City3D on PoznanRD. RoofDiffusion enhances gable details (Fig. 9), complex geometry recovery (Figs. 9 and 9), and is robust to tree noise (LABEL:fig:city3d-d).
@@ 6.4 Qualitative Evaluation of Height Completion
Figure 7: Evaluation of completion and denoising of real-world scans.
@@
Real-world Noise. We artificially remove points from the rasterized height maps to evaluate the resilience of our model to such corruptions, even in the presence of real-world noise and unseen roof geometries in both AHN3 [29] and Dales3D [82]. The 99% sparsity (Figs. 6(a) and 6(c)) and 70% incompleteness (Figs. 6(b) and 6(d)) are synthesized. We observe that IDW [74] produces a mosaic effect, particularly in incomplete regions, and fails to capture the true geometry of roofs. Meanwhile, Perona-Malik Diffusion [3] yields smoother outcomes but tends to overlook sharp roof details. In contrast, RoofDiffusion effectively yields clean and faithful geometric reconstructions.
Real-world Incompleteness. We collect point clouds exhibiting incompleteness from real-world scans, sourced from the Wayne County dataset [78]. Notably, the point cloud is rated at QL2 [80], which is the most common quality of LiDAR scans across the United States. Figures 6(e) and 6(f) show that RoofDiffusion effectively leverages available points in conjunction with footprint information to reconstruct missing parts, even when these parts are distant from the existing points. In contrast, IDW [74] and Perona-Malik Diffusion [3] failed in such scenarios due to a lack of prior information about roof structures. Figures 9 and 9 also shown No-FP RoofDiffusion can recover incompleteness with cleaner and sharper features.
Real-world Tree Noise. We also gather point clouds containing tree noise from the Cambridge dataset [79]. As illustrated in Figs. 6(g), 6(h) and 6(i), RoofDiffusion demonstrates its efficacy in restoring roof geometry precisely while eliminating the tree points. Moreover, Figs. 9 and 9 highlights that No-FP RoofDiffusion can effectively extract and recover roof geometry from overlapping tree noise.
Figure 8: 3D reconstruction using different height map pre-processors. RoofDiffusion matches the ground truth (GT) meshes generated by feeding GT height maps into City3D [29], except for (b) where RoofDiffusion omits a dormer present in GT, this discrepancy is reasonable due to the lack of data points in that area.
@@
Figure 8: 3D reconstruction using different height map pre-processors. RoofDiffusion matches the ground truth (GT) meshes generated by feeding GT height maps into City3D [29], except for (b) where RoofDiffusion omits a dormer present in GT, this discrepancy is reasonable due to the lack of data points in that area.
Figure 9: Evaluation of completion and denoising in real-world scans without footprints.
@@ 6.5 Limitations
We selected data containing tree noise from USGS 3DEP LiDAR sampled over Wayne County, MI [78], and Cambridge, MA [79], for analysis. In approximately 36% of these cases, tree noise was mistakenly reconstructed as building structures, see examples in the supplement. This error was due to the resemblance of tree canopies to architectural elements, such as dormers or chimneys, or due to severe occlusion by tree noise obscuring the underlying structures. Footprint-guided RoofDiffusion can encounter challenges due to severe misalignment between the footprint and the height map. This problem also occurs in [29]. Incorporating misalignment into the data augmentation pipeline and relaxing the footprint mask $\boldsymbol{m}$ during the denoising process warrants further investigation. Moreover, factors such as floor space, height from the ground, and location are useful indicators of roof geometry but are not yet used in the model. We leave the above as future work.
@@ 7 Conclusions
8 Acknowledgement
We thank Oleksandr Moskalenko and Dr. Zhe Jiang for generously providing computational resources on UF HyperGator. We also thank all Meta colleagues for helping in processing data and providing feedback throughout this project. Additionally, we acknowledge the funding provided by the UF Graduate School Preeminence Award and the UFII Fellowship.

### v2

@@ Abstract
Accurate completion and denoising of roof height maps are crucial to reconstructing high-quality 3D buildings. Repairing sparse points can enhance low-cost sensor use and reduce UAV flight overlap. RoofDiffusion is a new end-to-end self-supervised diffusion technique for robustly completing, in particular difficult, roof height maps. RoofDiffusion leverages widely-available curated footprints and can so handle up to 99% point sparsity and 80% roof area occlusion (regional incompleteness). A variant, No-FP RoofDiffusion, simultaneously predicts building footprints and heights. Both quantitatively outperform state-of-the-art unguided depth completion and representative inpainting methods for Digital Elevation Models (DEM), on both a roof-specific benchmark and the BuildingNet dataset. Qualitative assessments show the effectiveness of RoofDiffusion for datasets with real-world scans including AHN3, Dales3D, and USGS 3DEP LiDAR. Tested with the leading City3D algorithm, preprocessing height maps with RoofDiffusion noticeably improves 3D building reconstruction. RoofDiffusion is complemented by a new dataset of 13k complex roof geometries, focusing on long-tail issues in remote sensing; a novel simulation of tree occlusion; and a wide variety of large-area roof cut-outs for data augmentation and benchmarking. Code and dataset [^1] Created and released by the University of Florida: github.com/kylelo/RoofDiffusion
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Corrupted Point Cloud Corrupted Height Map Denoise 50% Denoise 90% Denoise 96% Restored Height Map Restored Point Cloud
@@
(a) Sparse
(b) Incomplete
(c) Tree Noise
@@ 1 Introduction
Thanks to the efforts of contributors worldwide, OpenStreetMap [56] now increasingly provides vectorized high-precision footprints [29, 24]. Consequently, recent building reconstruction algorithms [29, 24, 2] have incorporated footprints into the reconstruction process. However, these approaches still struggle with sparsity, incompleteness and noise. Footprint-guided RoofDiffusion addresses these long-tail challenges (Fig. 3).
@@
  • We unveil the PoznanRD (Poznan Roof Dataset) with 13k Level of Detail (LoD) 2.2 [4] noise-free roof meshes and height maps. When treated with our noise and incompleteness algorithm, our complex roofs dataset can effectively generate ample training data to address “long tail” challenges, see Fig. 3.
@@ 4.2 Height Completion based on a Diffusion Model
Given $\boldsymbol{x}$, our goal is to generate a predicted image, $\hat{\boldsymbol{x}}_{0}$, that closely approximates the normalized ground truth image, $\boldsymbol{x}_{\text{gt}}$. We conceptualize this height completion task as a diffusion process. Drawing on the work of [25], our approach employs both a forward and a reverse Markov chain. The forward chain perturbs the ground truth data through noise injection, while the reverse chain utilizes a learnable model, conditioned on the input, to restore the data, see Footnote 3.
@@
The Reverse Process sequentially removes noise from the data to yield a complete, noise-free version. Specifically, we can draw a noisy image $\boldsymbol{x}_{T}$ from $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ at time $t=T$. Then use a reverse transition kernel to recursively eliminate noise until $t=0$, resulting in the restored height image, $\hat{\boldsymbol{x}}_{0}$. This transition kernel is parameterized by a learnable model $\boldsymbol{\epsilon}_{\theta}$. While $\hat{\boldsymbol{x}}_{0}$ conforms to $q(\boldsymbol{x}_{\text{gt}})$, it may not always correspond to the normalized corrupted height image, $\boldsymbol{x}$. In our case, the objective is to ensure that the repaired roofs seamlessly integrate with pre-existing structures. To ensure a close association between $\boldsymbol{x}$ and $\hat{\boldsymbol{x}}_{0}$, we condition $\boldsymbol{\epsilon}_{\theta}$ on $\boldsymbol{x}_{t}$, $\boldsymbol{x}$, and $\bar{\alpha}_{t}$, in the spirit of [70]. Each step of the reverse process can be formulated as
@@
where $\alpha_{t}$ and $\bar{\alpha}_{t}$ are hyperparameters for variance scheduling, and $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
@@ 5 Datasets & Benchmarks
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 6: Sparsity of the real-world USGS 3DEP LiDAR dataset [79, 78].
@@ 5.1 Tree Points
In analyzing height maps within a footprint, tree crowns may intrude, causing LiDAR to capture the tree canopy instead of the roof, see Fig. 2(c). Thus, we simulate tree intrusion by planting virtual trees around the roof. We collected a real-world database with 1k tree height maps from [10]. Small trees are randomly placed outside the footprint, with random rotations and height adjustments. A “max” operation is used to combine the roof $\boldsymbol{x}$ and tree height map $\boldsymbol{x}_{\text{tree}}$. Algorithm 3 in supplements has the details.
@@ 6.1 Quantitative Evaluation of Height Completion
    (No-FP) Both sparsity and incompleteness in the PoznanRD in Tab. 2.
@@
(No-FP) Both sparsity and incompleteness. We compare to the state-of-the-art unguided depth completion methods, pNCNN [15] and CU-Net [85], that we selected for their exclusive use of depth data and hence close relation to height completion tasks. These algorithms were trained on our PoznanRD, using identical data augmentation settings, but no footprint information was used. We adhered to the default hyperparameter configurations specified for each method. To ensure a fair comparison, we also trained No-FP RoofDiffusion, a variant of RoofDiffusion that does not require footprints. We tested reconstruction accuracy on PoznanRD (s95 i30, s95 i60, s99 i30, and s99 i60) utilizing MAE and RMSE. The noise injection settings mirror the experiment in Tab. 1. Since predicting both footprint and roof simultaneously is a more difficult problem, the sparsity and incompleteness selection in Tab. 2 is lower than in Tab. 1.
@@
(No-FP) Either sparsity or incompleteness. When tested on the BuildingNet [73] dataset, No-FP RoofDiffusion outperforms linear interpolation, CU-Net [85], and pNCNN [15] in scenarios with dominant sparsity (s98 i0, s99 i0) or incompleteness (s90 i60, s90 i80). While CU-Net’s smoother predictions reduce large errors and improve RMSE, notably at s90 i80, s98 i0, and s99 i0 in Tab. 2, CU-Net does not predict sharp height maps well, as indicated by a higher MAE.
@@ 6.3 Enhancement of 3D Reconstruction
Figure 9 illustrates that using RoofDiffusion as a preprocessing step significantly enhances the reconstruction quality in City3D on PoznanRD. RoofDiffusion enhances gable details (Fig. 9), complex geometry recovery (Figs. 9 and 9), and is robust to tree noise (Fig. 9).
@@ 6.4 Qualitative Evaluation of Height Completion
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i)
Figure 7: Evaluation of the completion and denoising of real-world scans.
@@
Real-world Noise. We artificially remove points from the rasterized height maps to evaluate the resilience of our model to such corruptions, even in the presence of real-world noise and unseen roof geometries in both AHN3 [29] and Dales3D [82]. The 99% sparsity (Figs. 7(a) and 7(c)) and 70% incompleteness (Figs. 7(b) and 7(d)) are synthesized. We observe that IDW [74] produces a mosaic effect, particularly in incomplete regions, and fails to capture the true geometry of roofs. Meanwhile, Perona-Malik Diffusion [3] yields smoother outcomes but tends to overlook sharp roof details. In contrast, RoofDiffusion effectively yields clean and faithful geometric reconstructions.
Real-world Incompleteness. We collect point clouds exhibiting incompleteness from real-world scans, sourced from the Wayne County dataset [78]. Notably, the point cloud is rated at QL2 [80], which is the most common quality of LiDAR scans across the United States. Figures 7(e) and 7(f) show that RoofDiffusion effectively leverages available points in conjunction with footprint information to reconstruct missing parts, even when these parts are distant from the existing points. In contrast, IDW [74] and Perona-Malik Diffusion [3] failed in such scenarios due to a lack of prior information about roof structures. Figures 9 and 9 also shown No-FP RoofDiffusion can recover incompleteness with cleaner and sharper features.
Real-world Tree Noise. We also gather point clouds containing tree noise from the Cambridge dataset [79]. As illustrated in Figs. 7(g), 7(h) and 7(i), RoofDiffusion demonstrates its efficacy in restoring roof geometry precisely while eliminating the tree points. Moreover, Figs. 9 and 9 highlight that No-FP RoofDiffusion can effectively extract and recover roof geometry from overlapping tree noise.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: 3D reconstruction using different height map pre-processors. RoofDiffusion matches the ground truth (GT) meshes generated by feeding GT height maps into City3D [29], except for (b). In (b) RoofDiffusion omits a dormer present in GT due to the lack of data points in that area.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Evaluation of the completion and denoising in real-world scans without footprints.
@@ 6.5 Limitations
We selected data containing tree noise from USGS 3DEP LiDAR sampled over Wayne County, MI [78], and Cambridge, MA [79], for analysis. In approximately 36% of these cases, tree noise was mistakenly reconstructed as building structures, see examples in the supplement. This error was due to the resemblance of tree canopies to architectural elements, such as dormers or chimneys, or due to severe occlusion by tree noise obscuring the underlying structures. Footprint-guided RoofDiffusion can encounter challenges due to severe misalignment between the footprint and the height map. This problem also occurs in [29]. Incorporating misalignment into the data augmentation pipeline and relaxing the footprint mask, $\boldsymbol{m}$, during the denoising process warrants further investigation. Moreover, factors such as floor space, height from the ground, and location are useful indicators of roof geometry but are not yet used in the model. We leave the above as future work.
@@ 7 Conclusions
Acknowledgement
We thank Oleksandr Moskalenko and Dr. Zhe Jiang for generously providing computational resources on UF HyperGator. We also extend our appreciation to our colleagues at Meta Platforms for their assistance in processing data and offering valuable feedback throughout this project. We acknowledge the map services and data available from the U.S. Geological Survey, National Geospatial Program, and the Spatial Information System of the City of Poznań, Poland. Additionally, we are grateful for the funding provided by the UF Graduate School Preeminence Award and the UF Informatics Institute Fellowship.